OpenStudy (anonymous):

Two particles move according to x1 = A + Bt^2 and x2 = Dt^3, t ≥ 0 , where A, B, and D are positive constants. At t = 0 the particles are a distance A apart, and as t increases they move apart, but there is a time t > 0 at which the particles are as close as they will be ever again; in other words, |x1 − x2| is as small as it will get. What is this smallest value of |x1 − x2| for t > 0?

OpenStudy (anonymous):

Seems we should define s = x1 -x2 then set ds/dt = 0 to find max or min and even check (d/dt)(ds/dt) to see whether we have a max or min. ds/dt = 2 B t - 3 D t^2 =0 which can occur at t=0 and when 2 B = 3 D t t= (2B/3D) (d/dt)(ds/dt) = 2B - 6 D t = 2B - 6 D (2B/3D) = 2B-4B = -2B <0 unfortunately, that indicates we have a maximum, which is hard to believe, as t^3 will dominate as t gets very large. Puzzling.

OpenStudy (anonymous):

1. A again 2. A +4/27 (B^3/D^2) 3. A +1/2 (B^3/D^2) 4. A + 5/42 (B^3/D^2) 5. A + 1/6 (B^3/D^2) 6. A +(B^3/D^2) 7. (A)( B)/D 8. A + 2/27 (B^3/D^2) 9. Zero

OpenStudy (anonymous):

i thought it ws A I still dont understand

OpenStudy (anonymous):

I guess that the two curves, x1 and x2, cross at some t>0, which would mean that the minimum separation is 0.